how many grams of CuO is produced if 1.10 g of copper metal is oxidized law of definite proportions


1
Expert's answer
2021-03-03T06:30:32-0500

2Cu + O2 = 2CuO

n = m / M

M (Cu) = 63.55 g/mol

M (CuO) = 79.55 g/mol

n (Cu) = 1.10 / 63.55 = 0.02 mol

n (CuO) = n (Cu)

m (CuO) = n x M = 0.02 x 79.55 = 1.60 g